Your client just turned 75 years old and plans on retiring in 10 years on her 85th birthday. She is saving money today for he$12.401.59 | $12.998.63 | $13,243.18 $13,759.44 | $14,021.53
Answer #1

Amount required on 85th birthday towards yearly payments during retirement is the present value of a growing annuity due, arrived at $ 259,432.32 as follows:

A B C D 1 Present Value of Annuity Payments at the beginningof each period 2 3 Present value of growing annuity is calculated

Value as on the 85th birthday, of the amount intended for descendants to be paid on 90th birthday is the discounted value of $200,000 arrived at $ 136,116.64 as follows:

1 A B C D E F 3 Discounted value of an amount is calculated using the formula V=F/(1+r)^n 4 Where F= Amount after the specifi

Therefore, total amount required as on 85th birthday= $259,432.32 +$136,116.64. = $395,548.96

Value of the current savings of $100,000 as on 85th birthday is the future value arrived at $215,892.50 as follows:

B C D G H A 1 Future Value 3 Future value of an amount is calculated using the formula FV=P(1+r)^n 4 Where P= Principal (Pres

Therefore, value of net amount to be saved as 85th birthday= $3995,548.96- $215,892.50= $179,656.46

Amount required to be saved every year till retirement is the annuity, arrived at $12,401.59 as follows:

Ε A B C D 4 Where PMT= Periodical payments required, 5 PV= Present value of future sum which is FV/(1+r)^n 6 n= Number of pay

Hence the answer is the first option give, ie., $12,401.59